OpenStudy (anonymous):

if log50=1.699, then what is the value of log100(50)

OpenStudy (anonymous):

3.699

OpenStudy (anonymous):

10^2=100

OpenStudy (anonymous):

log100=2

OpenStudy (anonymous):

log100(50)=log100+log50=2+1.699=3.699

OpenStudy (anonymous):

these are my answer choices ... a.0.01699 b. 0.849 c. 169.9 d. 0.0849

OpenStudy (anonymous):

@Jack119

OpenStudy (anonymous):

is it log(100^(50)) or log(100(50))

OpenStudy (anonymous):

the question is not clear what are the bases?

OpenStudy (anonymous):

You can assume its base 10, sense 10^{1.699} is very close to 50

OpenStudy (anonymous):

the tex isn't formatted properly tho

OpenStudy (anonymous):

my guess is the question is asking this: if \[\log_{10}(50)=1.699\] then what is \[\log_{100}(50)\]

OpenStudy (anonymous):

we can solve with the change of base formula

OpenStudy (anonymous):

the first one is Log (base 50)=1.699 then it asks for the value of log(subscript 100 (50)

OpenStudy (anonymous):

\[\log_{100}(50)=\frac{\log(50)}{\log(100)}=\frac{\log(50)}{2}=\frac{1.699}{2}\]

OpenStudy (anonymous):

in plain english, it is half of \(1.699\)

OpenStudy (anonymous):

oh okay so basically divide the 1.699 by two and thats my answer correct?